OpenStudy (anonymous):

Which is equivalent to (5z + 15)^(4) ? a. 625(z + 15)^(4) b. 625(z + 3)^(4) c. 5(z + 15)^(4) d. 5(z + 3)^(4) explanation too please

OpenStudy (mertsj):

\[5z+15=5(z+3) so (5z+15)^4=[5(z+3)]^4=5^4(z+3)^4=625(z+3)^4\]

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw